SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)


Page no longer updated

Status of DIRC Electronics, PMT's and HV System

Last updated: 06-01-02 (Muriel) 


Although all related data should soon be available in the database, this page provides information about the current status, the configuration and quality control of the Dirc Front-End  Boards (DFB), the PMT's and related hardware.

 

Front-End Boards:

  • Status: All dfb's are operational. Reprogramming of the seqdec alteras using the newest version of <seqdec-02-99.jam> failed due to communication problems between the boards and dataflow. The previous program version <seqdec-12-98.jam> was restored. The DCC's were reprogrammed in Feb. 1999 by C. Renard in order to remove the calibration strobe dependence on the address tag sent with each opcode. Two RadFETs were installed on sector 3+8 (03-03-99) to continously monitor the background level. They are connected to channels 28 of the west and east-side EMC radiation monitor boards. There was once again an occurence of the out-of-order problem in sector 7 (DFB 0x42). It was removed by exchanging the DFB (08-03-99).
  • Configuration: The configuration file used for the DFB DAC gains is: <dfbsetup-98.txt>. The physical location of the DFB's in the FEE crates with respect to their serial board numbers is retrieved by a dedicated "calibration" cycle. An actualized version of Gerard's <drivingDirc-03-05-99.ps> is now available.
  • Calibration: The current electronics gain and pedestal calibration file, needed to transform measured counts into mVolts: <dfbcalib-07-98.txt>
  • Problems with boards: The following DFB's have shown failures during running or testing modes (start of list: August 1999):
DFB SN Sector Harness Comment Date of discovery/understanding
42 07 12 Bad trigger tag FIFO (out-of-order damages) 08-03-99
?? 07 09 Bad voltage readout  08-03-99
?? 05 01 Bad analog chip  07-01-99

 

Photomultiplier Tubes:

  • Status: LED and noise runs have shown that about 0.2% of the PMT's are noisy or inefficient (Oct 99). During the October shutdown, the inspection of the tubes revealed 50 crazed ("frosty") tubes. The list of these, and some specially milky tubes is found at < frostyPMT.html>. In March 2002, the 6.4 million DIRC standalone random triggers allowed to look at dead/inefficient/noisy PMTs with high statistics. New ones were then discovered (Jochen). This figure and the following table lists the location of these tubes:

Dead tubes   (zero Hz):            
                 
sector harness channel column row rate (Hz) number problem discovery
2 8 63 18 40 0 1 ?? 3/10/02
4 0 0 0 7 0 2 ?? 3/10/02
7 11 60 22 39 0 3 masked 2/25/02
7 11 61 23 39 0 4 masked 3/1/02
9 13 42 27 32 0 5 masked(X-tree) 8/8/01
Very inefficient tubes ( < 1Hz):          
                 
sector harness channel column row rate (Hz) number problem discovery
0 3 39 6 30 0.52 1 ?? 10-Mar
(old)   0 3 46 8 34 0.26 2 disconnected 4/26/00
(old)   1 3 7 3 7 0.52 3 dead ? Feb-99
2 2 46 5 33 0.52 4 disconnected 8/1/01
2 10 30 16 19 0.52 5 disconnected 1/3/02
3 5 31 10 25 0.79 6 ?? 3/10/02
7 2 25 3 21 0.79 7 disconnected 1/3/02
7 2 29 3 24 0.52 8 disconnected(X-tree) 1/3/02
10 2 47 4 34 0.52 9 ?? 3/10/02
Inefficient   tubes (1-30 Hz):            
                 
sector harness channel column row rate (Hz) number problem discovery
0 0 30 0 23 2.1 1 disconnected 8/1/01
0 3 7 3 7 4.71 2 ?? 3/10/02
(old)   0 7 62 15 40 1.31 3 dead ?old?
1 1 45 1 33 5.5 4 ?? 3/10/02
1 12 1 16 9 9.17 5 ?? 3/10/02
1 12 52 25 36 1.31 6 dead 1/27/01
(old)   2 0 24 0 20 8.38 7 dead-removed 5/24/00
2 0 54 1 35 2.88 8 ?? 3/10/02
(old)   2 1 4 0 4 4.98 9 disconnected 4/26/00
(old)   2 1 12 2 10 3.4 10 disconnected 4/26/00
2 10 8 12 5 4.71 11 ?? 3/10/02
(old)   2 13 63 29 40 1.57 12 used to determine machine bkg 4/1/99
3 1 6 1 5 3.4 13 ?? 3/10/02
3 1 57 3 38 11 14 ?? 3/10/02
(old)   3 2 26 4 21 2.36 15 ?? ?
(old)   3 2 29 3 24 1.83 16 ?? ?
(old)   3 2 61 6 39 3.14 17 dead-removed 5/28/00
3 5 61 12 39 3.4 18 ?? 3/10/02
3 7 30 10 19 2.1 19 ?? 3/10/02
3 10 23 15 15 25.67 20 ?? 3/10/02
3 13 40 25 32 23.31 21 ?? 3/10/02
4 0 27 1 21 1.05 22 disconnected 4/4/01
(old)   4 2 1 1 1 2.88 23 dead-noisy-removed 5/1/99
(old)   4 13 14 21 23 2.62 24 disconnected 4/26/00
(old)   5 4 19 5 19 2.88 25 dead 2/1/99
5 8 10 10 6 20.69 26 ?? 3/10/02
5 8 23 12 15 5.24 27 ?? 3/10/02
5 12 0 15 8 6.02 28 ?? 3/10/02
(old)   5 13 2 18 15 6.29 29 disconnected(X-tree) 4/26/00
(old)   5 13 15 22 23 1.57 30 ??? ?/00
6 0 37 1 26 6.29 31 ?? 3/10/02
6 1 27 2 22 2.36 32 ?? 3/10/02
(old)   6 5 3 4 3 2.1 33 dead 5/1/99
6 6 55 14 35 7.33 34 ?? 3/10/02
6 7 13 8 9 1.05 35 ?? 3/10/02
6 9 17 12 11 3.14 36 ?? 3/10/02
6 9 28 15 18 24.36 37 ?? 3/10/02
6 10 3 11 2 2.1 38 ?? 3/10/02
7 1 42 3 31 9.43 39 ?? 3/10/02
7 2 24 3 20 7.33 40 ?? 3/10/02
7 7 1 11 15 24.1 41 ?? 3/10/02
(old)   7 8 24 11 15 2.1 42 removed-X-tree 8/24/00
(old)   7 11 51 20 34 5.5 43 ??? ?/00
(old)   8 0 53 0 35 8.9 44 dead ?/00
8 0 54 1 35 8.9 45 ?? 3/10/02
8 2 54 6 37 6.55 46 ?? 3/10/02
(old)   8 3 2 3 2 3.93 47 noisy(100kHz_ 6/1/00
(old)   8 3 46 8 34 2.88 48 dead ?/00
8 7 63 16 40 20.95 49 ?? 3/10/02
8 8 11 9 7 3.14 50 ?? 3/10/02
8 8 32 13 21 4.45 51 ?? 3/10/02
8 8 47 16 31 18.6 52 ?? 3/10/02
8 9 16 11 11 6.02 53 X-tree 3/10/02
8 13 11 22 21 2.1 54 ?? 3/10/02
(old)   8 13 63 29 40 6.02 55 dead ?/00
(old)   9 3 22 5 20 2.1 56 dead ?/00
9 8 6 8 4 3.93 57 ?? 3/10/02
9 8 9 9 6 8.12 58 ?? 3/10/02
9 9 53 18 35 1.57 59 ?? 3/10/02
9 12 2 16 10 7.07 60 ?? 3/10/02
(old)   9 12 20 19 23 4.45 61 dead ?/00
10 0 6 0 11 2.62 62 ?? 3/10/02
10 1 4 0 4 7.6 63 ?? 3/10/02
10 6 29 10 20 2.1 64 ?? 3/10/02
11 7 36 13 22 1.83 65 ?? 3/10/02
(old) 11 8 42 15 27 1.31 66 ??? ?/00
11 10 7 12 4 4.71 67 ?? 3/10/02
Noisy tubes ( > 1kHz):            
                 
sector harness channel column row rate (Hz) number problem discovery
0 12 52 25 36 6776 1 ?? 3/10/02
1 3 38 5 30 1229 2 ?? 3/10/02
2 0 8 0 12 5070 3 ?? 3/10/02
6 1 34 3 26 6084 4 ?? 3/10/02
6 2 12 3 8 1119 5 ?? 3/10/02
6 5 42 10 30 1856 6 ?? 3/10/02
8 12 3 17 10 1307 7 disconnected 4/4/01
9 3 37 7 29 1483 8 ?? 3/10/02
9 7 39 13 24 4500 9 ?? 3/10/02
9 10 3 11 2   10   11/5/01
10 2 32 4 26 1590 11 ?? 3/10/02
10 8 43 15 28 2878 12 ?? 3/10/02
  • Status (2): Some PMT's showed a milky surface of presently unknown consistency .he following list is far from being complete.
Sector Harness Channel Column Row degree of milky surface
8 5 22 7 20 +++
8 3 5 3 5 +++
8 4 9 5 9 ++
5 5 8 6 8 +
  • Changes: Any exchange of badly working PMT's has been postponed to the September '99 SOB opening
  • Cable errors_1: (03/29/99): A preliminary analysis using cosmic data reveals about 12 candidates for miscabled PMT signal wires. More information is available on a dedicated web page: <cableerrors-03-15-99.html>.
  • Cable errors_2: (04/11/00): A new study was perfomed, by Joe Schwiening, using high statistic reconstructed thetaC data. More information is available on this page: <cableerrors-04-11-00.html>.

 

Light Catchers (LC):

  • Status: After the opening of the SOB (1st of March '99), the LC look pretty well. An analysis of LED data (Feb-99), comparing the rows without LC (sector 11, rows 24-26)  to same rows in other sectors, proved a gain on the photon detection efficiency of (20 +/- 2)%.
  • Changes: The missing 3 rows of LC were mounted during the SOB opening the 03-05-99. 1 LC panel covering 10 PMT's in a less occupied region of sector 6 (see table below) was removed in order to be able to control quality and aging of the LC.
Sector Harness Channel Column Row Status 
11 0-13 32-36 0-25 24-26 mounted the 03-05-99
6 0-2 59-63 0-4 39-40 removed the 03-05-99
 

High Voltage CAEN crates:

  • Status: The door closing after the October shutdown and opening of the SOB doors has damaged two HV pipes carrying each 60 HV cables and supplying sector 0 and ? (4 spares per sector). Four cables in sector 0 are known to be broken. Currently, this affects group 31 and 33, which are set to low voltage (about 30 V) in order to avoid alarms.
  • Configuration: HV configuration file (FORTRAN): <hvsetup-98.cc>
  • Miscellaneous: The HV control VME board (IOC) can be reached using the command: xyplex drc-hv. All SY527 crates may be directly addressed via a serial link and the xyplex terminal server using: xyplex drc-caeni, with i=1,...,6
  • Problems: Problems with particular HV channels (groups) are listed below:
Sector Harnesses Channels HV-groups Comments Date of discovery
00 00-00 00-00 00 00. 
00-00-00
  • Changes: The following table gives the exchanged HV groups during or after the cosmic run ("<>" is "exchanged"):
Sector Harnesses Channels HV-groups Volts Date of exchange
0 03 07 <> 09  04 <> 09 1107 <> 1054 02-23-99
5 06,07 <> 02,03 08-15 <> 08-15 08 <> 09 1130 <> 0980 02-23-99
6 12,13 <> 10,11 08-15 <> 24-31 18 <> 19 1124 <> 1073 02-23-99
6 06 <> 05 05 <> 06 02 <> 03 1318 <> 1228 02-23-99
8 02,03 <> 00,01  00-07 <> 00-07 04 <> 05 1326 <> 1135 02-23-99
8 12,13 <> 00,01 40-47 <> 48-55 44 <> 45 1350 <> 1193 02-23-99
11 12,13 <> 04,05 32-39 <> 40-47 38 <> 39 1385 <> 1099 02-03-99

 

BaBar DIRC Home Page
DIRC Operations Page


This page is maintained by George Vasileiadis